# Anomaly-induced vanishing of brane partition functions

### v1

@@ 5 Chern-Simons theories
for some integer level $\kappa$. The dimension of the Hilbert space of the chiral gauge theory scales as some power of $|\kappa|$ Hsieh:2020jpj . If we assume that the partition function is a section of a line bundle, then we shall hereafter restrict our attention to invertible theories where $\kappa=\pm 1$, corresponding respectively to self-dual and anti-self-dual gauge fields.
Note that such an anomaly is not well-defined as an element of $\mathbb{R}/\mathbb{Z}$ when $\kappa$ is odd, so a precise formulation of it requires the introduction of a quadratic refinement $q(\underline{\check{C}})$ Hopkins:2002rd , [^25] Unlike the quadratic refinements we used in Section 4 which are defined for intersection pairings, those introduced here are for torsion pairings. defined to be such that
@@
To simplify notation, it is customary to define another quadratic refinement $\tilde{q}(\underline{\check{C}})\coloneqq q(\underline{\check{C}})-q(0)$. Following Hsieh:2020jpj , the anomaly can be expressed as
@@
where $\mathcal{A}_{\text{grav}}$ is the gravitational anomaly of the theory. [^26] The gravitational anomaly is given by $\mathcal{A}_{\text{grav}}=\eta(\mathcal{D}_{Y_{3}}^{\text{Dirac}})$ for $d=2$, and $\mathcal{A}_{\text{grav}}=28\eta(\mathcal{D}_{Y_{7}}^{\text{Dirac}})$ for $d=6$, where $\mathcal{D}_{Y_{d+1}}^{\text{Dirac}}$ is the Dirac operator on $Y^{d+1}$, and $\eta$ is the corresponding eta-invariant Hsieh:2020jpj . For $d=2,6$, if $Z^{d+2}$ admits a Spin structure, then one can further pick the convention $q(0)=-\mathcal{A}_{\text{grav}}$ such that $\mathcal{A}[\underline{\check{C}}]=-\kappa q(\underline{\check{C}})$, with
@@ 5.1 Integral lifts of Wu classes
where $\underline{\Lambda}_{p+2}$ is defined such that its characteristic class $\underline{\Lambda}_{p+2}]_{\mathbb{Z}}\in H^{p+2}(Z^{d+2};\mathbb{Z})$ is an integral lift of the $(p+2)$-th Wu class $v_{p+2}\in H^{p+2}(Z^{d+2};\mathbb{Z}_{2})$ Witten:1996md ; Hopkins:2002rd . To see why, note that on an $n$-manifold $M^{n}$, the Wu class $v_{i}$ is defined to be a class representing the Steenrod square operation $\text{Sq}^{i}:H^{\ast}(M^{n};\mathbb{Z}_{2})\to H^{\ast+i}(M^{n};\mathbb{Z}_{2})$ 10.2307/j.ctt1b7x52h.1 , such that
@@
for any $x_{n-i}\in H^{n-i}(M^{n};\mathbb{Z}_{2})$ 5895d0ff-dbfd-39a7-9ffe-3ef746a10f70 . We also have that, by definition, $\text{Sq}^{j}(x_{j})=x_{j}\cup x_{j}$ for any $j$. For $n$ even, if we take $i=n/2$, then
@@
On spin manifolds, $w_{1}=w_{2}=0$, so the second Wu class $v_{2}$ vanishes identically. Its integral lift can be chosen to be trivial, as we saw in (132) for the case of $d=2$ and $p=0$. Similar remarks apply to the case of $d=4$ and $p=1$. When $d=6$ and $p=2$, the fourth Wu class reduces to $w_{4}$, then the relation 3dcb6f87-ee89-3c37-b8a3-6317f6f6a26b
@@
Let us start with the chiral boson in $d=2$. Due to degree reasons, any 2-manifold $X^{2}$ must have $V_{3}=W_{3}=0$, i.e. it admits a $\text{Pin}^{c}$ structure. [^27] As a remark, a $\text{Pin}^{+}$ structure requires $w_{2}=0$, a $\text{Pin}^{-}$ structure requires $w_{2}+w_{1}^{2}=0$, and a Spin structure requires $w_{1}=w_{2}=0$. Similarly, a $\text{Pin}^{c}\coloneqq\text{Pin}^{\pm}\times_{\mathbb{Z}_{2}}\mathrm{U}(1)$ structure requires $W_{3}=0$, while a $\text{Spin}^{c}\coloneqq\text{Spin}\times_{\mathbb{Z}_{2}}\mathrm{U}(1)$ structure requires both $w_{1}=0$ and $W_{3}=0$. The integral lift of $v_{2}=w_{2}+w_{1}^{2}$ is given by
@@
On top of that, there is a contribution from the second integral Stiefel-Whitney class $W_{2}$, whose mod 2 reduction is $w_{1}^{2}$. Similarly to the $\text{Pin}^{c}$ structure whose obstruction is measured by $W_{3}$, we will refer to an $\text{SO}^{c}$ structure as the tangential structure whose obstruction is measured by $W_{2}$. This can be seen via a diagram analogous to (146) as follows.
@@
As we can infer above, the orthogonal structure of a manifold can be lifted to an $\text{SO}^{c}$ structure if and only if $W_{2}=0$, in which case commutativity of the diagram implies $w_{1}=f_{1}$ mod 2, where $f_{1}$ can be interpreted as the characteristic class of the axionic 1-form field strength of the compact scalar. This is an analogue of our earlier Maxwell example in Section 3.2, wherein the $\text{Spin}^{c}$ structure implies that $w_{2}$ coincides mod 2 with $F_{2}$, which is the field strength of the dynamical gauge field.
As an aside, we can similarly interpret the vanishing of $V_{i+1}=\beta(v_{i})$ as giving rise to a $\text{Wu}_{i}^{c}$ structure Sati:2011rw , which is depicted by the diagram below.
@@
A class of examples of manifolds which does not admit an $\text{SO}^{c}$ structure, i.e. $W_{2}\neq 0$, is those admitting a $\text{Pin}^{-}$ structure but not a $\text{Pin}^{+}$ structure. Recall that all 2-manifolds admit a $\text{Pin}^{-}$ structure but not necessarily a $\text{Pin}^{+}$ structure Kirby_Taylor_1991 . This implies all orientable 2-manifolds admit a Spin structure (and of course an $\text{SO}^{c}$ structure). For non-orientable 2-manifolds, a simple example that does not admit an $\text{SO}^{c}$ structure is $\mathbb{RP}^{2}$, which has $w_{2}=w_{1}^{2}\neq 0$.
@@ 5.3 M5-brane worldvolume theory
For the M5-brane, its anomaly theory can be modeled by a Chern-Simons theory in $d=6$ with the M-theory 3-form $C_{3}$ Witten:1996hc . The 11d M-theory spacetime is often assumed to admit a Spin structure, and typically one considers the wrapping of an M5-brane on a spin manifold, in which case (132) does the job. Particularly, $\Lambda_{4}=\frac{1}{2}p_{1}$ is a suitable integral lift of $v_{4}$. To keep the subsequent discussion general, we do not assume any tangential structure on the worldvolume $X^{6}$ of the M5-brane.
Note that Wu classes in degrees above half the dimension of the manifold vanish, so $v_{4}=0$ on 6-manifolds and thus $\Lambda_{4}$ exists. [^29] This follows from the axiom that $\text{Sq}^{i}(x_{j})=0$ if $i>j$ 5895d0ff-dbfd-39a7-9ffe-3ef746a10f70 . In addition, recall from (142) that the terms in $v_{4}$ which do not automatically admit integral lifts are $w_{4}$ and $w_{2}^{2}$, but $V_{5}=\beta(v_{4})=0$ on $X^{6}$ implies that the combination $w_{4}+w_{2}^{2}$ must be a mod 2 reduction. [^30] Even though $\text{Sq}^{1}(w_{2}^{2})=0$, the quantity $\beta(w_{2}^{2})$ is generally a non-vanishing even integral class. In contrast, $\beta(w_{1}^{2})=\beta(W_{2}\ \text{mod}\ 2)$ must be vanishing as an integral class by exactness. Let us denote it as $J_{4}$. In short, we can parametrise the integral lift of $v_{4}$ as
@@
It was argued by Diaconescu:2003bm that the partition function being non-vanishing is necessary for the M5-brane to decouple from the dynamics of the 11d supergravity bulk. Only when satisfied can the M5-brane worldvolume theory be treated as a standalone 6d (2,0) superconformal field theory (SCFT) Witten:1995em .
@@
Lastly, if $X^{6}$ is spin, then $w_{4}=\frac{1}{2}p_{1}$ mod 2 and $W_{4}=w_{2}=W_{2}=0$, so we obtain [^31] As was shown in Hsieh:2020jpj , $\frac{1}{4}p_{1}$ is indeed always an integral cohomology class on spin manifolds of dimension less than or equal to 7, due to the fact that $v_{4}=w_{3}=w_{2}=w_{1}=0$. Similar arguments apply to the previous cases.
@@
Our result for the spin case is indeed compatible with the well-known shifted quantisation of the $G_{4}$ flux in M-theory Witten:1996md , where
@@
The Freed-Witten anomaly cancellation condition (82) can be interpreted as the D3-brane admitting a twisted $\text{Spin}^{c}$ structure Freed:1999vc , with the twist given by $\mathsf{H}_{3}]_{\mathbb{Z}}$ or $\mathsf{G}_{3}]_{\mathbb{Z}}$. There is a similar interpretation for (165). Note that the quantity $\frac{1}{2}p_{1}$ is also known as the first fractional Pontryagin class. It measures the obstruction to lift a Spin structure to a String structure Killingback:1986rd ; Stolz_Teichner_2004 , as depicted by the diagram below.
@@
Therefore, we can interpret (165) as the M5-brane admitting a twisted String structure Sati:2009ic ; Sati:2010dc , with the twist given by $\mathsf{G}_{4}+\frac{1}{4}\mathsf{p}_{1}]_{\mathbb{Z}}$.
On a related note, based on the arguments in Witten:1998xy ; Freed:1999vc leading to (82), it was conjectured by Witten:1999vg that the analogue for the M5-brane should take the form,
@@
For completeness, we expect that (158) should be modified in the presence of sources, analogously to (83). By M/F-theory duality Johnson:1997wf , the Wilson operators are now M2-branes ending on a 2-submanifold of $X^{6}$. Schematically, the non-vanishing condition for the resulting M5-brane partition function becomes
@@
As alluded to earlier, this relation should be interpreted with care since the M5-brane can no longer be regarded as a decoupled theory with respect to the M-theory background Diaconescu:2003bm .
@@ 6 D3-brane in F-theory backgrounds
which is equivalent to $H_{\ast}(T^{2};\mathbb{Z})$ regarded as a $\mathbb{Z}[\pi_{1}(X^{4})]$-module b435f69a-14cc-3b23-b856-e21562c71558 ; MR1867354 .
Notably, in the middle degree we have $\text{Aut}_{\mathbb{Z}}(\mathbb{Z}\oplus\mathbb{Z})=\text{GL}(2,\mathbb{Z})$, i.e. the mapping class group of the torus. For our purposes, it suffices to restrict to the orientation-preserving subgroup $\text{SL}(2,\mathbb{Z})$. As a result, we interpret the resultant D3-brane to be placed in a Type IIB string theory background admitting a non-trivial $\text{SL}(2,\mathbb{Z})$-action. One familiar example of such is an orientifold background, e.g. $X^{4}=S^{1}\times\mathbb{RP}^{3}\subset\text{AdS}_{5}\times\mathbb{RP}^{5}$, where $\rho$ induces a flip in sign of the $\text{SL}(2,\mathbb{Z})$-doublet $([\mathsf{H}_{3}],[\mathsf{G}_{3}])\in H^{3}(\mathbb{RP}^{3};(\mathbb{Z}\oplus\mathbb{Z})_{\rho})$ as we go around a non-contractible loop Witten:1998xy . [^32] Whenever the context is clear, we will abbreviate $(\mathbb{Z}\oplus\mathbb{Z})_{\rho}$ as $\mathbb{Z}^{2}_{\rho}$. Using the M5-brane basepoint anomaly (157) as the starting point, our goal is to derive the Freed-Witten anomaly cancellation condition for the D3-brane in generic F-theory backgrounds by reducing over the fiber $T^{2}$. [^33] As a reminder, this torus $T^{2}$ is independent from the mapping torus $X^{d}\times S^{1}$.

### v3

@@ 5 Chern-Simons theories
for some integer level $\kappa$. The dimension of the Hilbert space of the chiral gauge theory scales as some power of $|\kappa|$ Hsieh et al. (2022). If we assume that the partition function is a section of a line bundle, then we shall hereafter restrict our attention to invertible theories where $\kappa=\pm 1$, corresponding respectively to self-dual and anti-self-dual gauge fields.
Note that such an anomaly is not well-defined as an element of $\mathbb{R}/\mathbb{Z}$ when $\kappa$ is odd, so a precise formulation of it requires the introduction of a quadratic refinement $q(\underline{\check{C}})$ Hopkins and Singer (2005), [^25] Unlike the quadratic refinements we used in Section 4 which are defined for intersection pairings, those introduced here are for torsion pairings. defined to be such that
@@
To simplify notation, it is customary to define another quadratic refinement $\tilde{q}(\underline{\check{C}})\coloneqq q(\underline{\check{C}})-q(0)$. Following Hsieh et al. (2022), the anomaly can be expressed as
@@
where $\mathcal{A}_{\text{grav}}$ is the gravitational anomaly of the theory. [^26] The gravitational anomaly is given by $\mathcal{A}_{\text{grav}}=\eta(\mathcal{D}_{Y_{3}}^{\text{Dirac}})$ for $d=2$, and $\mathcal{A}_{\text{grav}}=28\eta(\mathcal{D}_{Y_{7}}^{\text{Dirac}})$ for $d=6$, where $\mathcal{D}_{Y_{d+1}}^{\text{Dirac}}$ is the Dirac operator on $Y^{d+1}$, and $\eta$ is the corresponding eta-invariant Hsieh et al. (2022). For $d=2,6$, if $Z^{d+2}$ admits a Spin structure, then one can further pick the convention $q(0)=-\mathcal{A}_{\text{grav}}$ such that $\mathcal{A}[\underline{\check{C}}]=-\kappa q(\underline{\check{C}})$, with
@@ 5.1 Integral lifts of Wu classes
where $\underline{\Lambda}_{p+2}$ is defined such that its characteristic class $\underline{\Lambda}_{p+2}]_{\mathbb{Z}}\in H^{p+2}(Z^{d+2};\mathbb{Z})$ is an integral lift of the $(p+2)$-th Wu class $v_{p+2}\in H^{p+2}(Z^{d+2};\mathbb{Z}_{2})$ Witten (1997b); Hopkins and Singer (2005). To see why, note that on an $n$-manifold $M^{n}$, the Wu class $v_{i}$ is defined to be a class representing the Steenrod square operation $\text{Sq}^{i}:H^{\ast}(M^{n};\mathbb{Z}_{2})\to H^{\ast+i}(M^{n};\mathbb{Z}_{2})$ Steenrod and Epstein (1962), such that
@@
for any $x_{n-i}\in H^{n-i}(M^{n};\mathbb{Z}_{2})$ Milnor and Stasheff (1974). We also have that, by definition, $\text{Sq}^{j}(x_{j})=x_{j}\cup x_{j}$ for any $j$. For $n$ even, if we take $i=n/2$, then
@@
On spin manifolds, $w_{1}=w_{2}=0$, so the second Wu class $v_{2}$ vanishes identically. Its integral lift can be chosen to be trivial, as we saw in (132) for the case of $d=2$ and $p=0$. Similar remarks apply to the case of $d=4$ and $p=1$. When $d=6$ and $p=2$, the fourth Wu class reduces to $w_{4}$, then the relation Thomas (1960)
@@
Let us start with the chiral boson in $d=2$. We view this as the gapless edge mode of the $\text{U}(1)_{1}$ Spin-CS theory, following Witten (1997a). Note that the chirality constraint $d\phi\stackrel{{\scriptstyle!}}{{=}}d\phi$ requires a choice of volume form and therefore only makes sense on orientable manifolds. However, as explained in Delmastro and Gomis (2021), the bulk theory has a time-reversal symmetry, such that it can be defined as well on non-orientable manifolds. For the purpose of this work, it suffices for us to analyze the bulk anomaly theory in its own right, assuming that it admits a gapless edge mode (as in Hsieh et al. (2022)) which we loosely refer to as the “chiral boson.” In fact, by invoking considerations of integral Wu structures, our goal is to precisely characterise the conditions under which such a boundary theory may become well-defined, and particularly, admit a non-vanishing partition function. We do not attempt to provide any Lagrangian description of the resultant theory.
Due to degree reasons, any 2-manifold $X^{2}$ must have $V_{3}=W_{3}=0$, i.e. it admits a $\text{Pin}^{c}$ structure. [^27] As a remark, a $\text{Pin}^{+}$ structure requires $w_{2}=0$, a $\text{Pin}^{-}$ structure requires $w_{2}+w_{1}^{2}=0$, and a Spin structure requires $w_{1}=w_{2}=0$. Similarly, a $\text{Pin}^{c}\coloneqq\text{Pin}^{\pm}\times_{\mathbb{Z}_{2}}\mathrm{U}(1)$ structure requires $W_{3}=0$, while a $\text{Spin}^{c}\coloneqq\text{Spin}\times_{\mathbb{Z}_{2}}\mathrm{U}(1)$ structure requires both $w_{1}=0$ and $W_{3}=0$. The integral lift of $v_{2}=w_{2}+w_{1}^{2}$ is given by
@@
On top of that, there is a contribution from the second integral Stiefel-Whitney class $W_{2}$, whose mod 2 reduction is $w_{1}^{2}$. Let us provide a physical interpretation of the case when such a class contributes non-trivially. Similarly to the $\text{Pin}^{c}$ structure whose obstruction is measured by $W_{3}$, we will refer to an $\text{SO}^{c}$ structure as the tangential structure whose obstruction is measured by $W_{2}$. This can be seen via a diagram analogous to (146) as follows.
@@
As we can infer from above, the orthogonal structure of a manifold can be lifted to an $\text{SO}^{c}$ structure if and only if $W_{2}=0$, in which case commutativity of the diagram implies $w_{1}=f_{1}$ mod 2, where $f_{1}$ can be identified as the characteristic class of the axionic 1-form field strength of the compact scalar. This is an analogue of our earlier Maxwell example in Section 3.2, wherein the $\text{Spin}^{c}$ structure implies that $w_{2}$ coincides mod 2 with $F_{2}$, which is the field strength of the dynamical gauge field.
As an aside, we can similarly interpret the vanishing of $V_{i+1}=\beta(v_{i})$ as giving rise to a $\text{Wu}_{i}^{c}$ structure Sati (2012), which is depicted by the diagram below.
@@
A class of examples of manifolds which does not admit an $\text{SO}^{c}$ structure, i.e. $W_{2}\neq 0$, is those admitting a $\text{Pin}^{-}$ structure but not a $\text{Pin}^{+}$ structure. Recall that all 2-manifolds admit a $\text{Pin}^{-}$ structure but not necessarily a $\text{Pin}^{+}$ structure Kirby and Taylor (1991). This implies all orientable 2-manifolds admit a Spin structure (and of course an $\text{SO}^{c}$ structure). For non-orientable 2-manifolds, a simple example that does not admit an $\text{SO}^{c}$ structure is $\mathbb{RP}^{2}$, which has $w_{2}=w_{1}^{2}\neq 0$.
@@ 5.3 M5-brane worldvolume theory
For the M5-brane, its anomaly theory can be modeled by a Chern-Simons theory in $d=6$ with the M-theory 3-form $C_{3}$ Witten (1997a). The 11d M-theory spacetime is often assumed to admit a Spin structure, and typically one considers the wrapping of an M5-brane on a spin manifold, in which case (132) does the job. Particularly, $\Lambda_{4}=\frac{1}{2}p_{1}$ is a suitable integral lift of $v_{4}$. To keep the subsequent discussion general, we do not assume any tangential structure on the worldvolume $X^{6}$ of the M5-brane.
Note that Wu classes in degrees above half the dimension of the manifold vanish, so $v_{4}=0$ on 6-manifolds and thus $\Lambda_{4}$ exists. [^29] This follows from the axiom that $\text{Sq}^{i}(x_{j})=0$ if $i>j$ Milnor and Stasheff (1974). In addition, recall from (142) that the terms in $v_{4}$ which do not automatically admit integral lifts are $w_{4}$ and $w_{2}^{2}$, but $V_{5}=\beta(v_{4})=0$ on $X^{6}$ implies that the combination $w_{4}+w_{2}^{2}$ must be a mod 2 reduction. [^30] Even though $\text{Sq}^{1}(w_{2}^{2})=0$, the quantity $\beta(w_{2}^{2})$ is generally a non-vanishing even integral class. In contrast, $\beta(w_{1}^{2})=\beta(W_{2}\ \text{mod}\ 2)$ must be vanishing as an integral class by exactness. Let us denote it as $J_{4}$. In short, we can parametrise the integral lift of $v_{4}$ as
@@
It was argued by Diaconescu et al. (2003b) that the partition function being non-vanishing is necessary for the M5-brane to decouple from the dynamics of the 11d supergravity bulk. Only when satisfied can the M5-brane worldvolume theory be treated as a standalone 6d (2,0) superconformal field theory (SCFT) Witten (1996).
@@
Lastly, if $X^{6}$ is spin, then $w_{4}=\frac{1}{2}p_{1}$ mod 2 and $W_{4}=w_{2}=W_{2}=0$, so we obtain [^31] As was shown in Hsieh et al. (2022), $\frac{1}{4}p_{1}$ is indeed always an integral cohomology class on spin manifolds of dimension less than or equal to 7, due to the fact that $v_{4}=w_{3}=w_{2}=w_{1}=0$. Similar arguments apply to the previous cases.
@@
Our result for the spin case is indeed compatible with the well-known shifted quantisation of the $G_{4}$ flux in M-theory Witten (1997b), where
@@
The Freed-Witten anomaly cancellation condition (82) can be interpreted as the D3-brane admitting a twisted $\text{Spin}^{c}$ structure Freed and Witten (1999), with the twist given by $\mathsf{H}_{3}]_{\mathbb{Z}}$ or $\mathsf{G}_{3}]_{\mathbb{Z}}$. There is a similar interpretation for (165). Note that the quantity $\frac{1}{2}p_{1}$ is also known as the first fractional Pontryagin class. It measures the obstruction to lift a Spin structure to a String structure Killingback (1987); Stolz and Teichner (2004), as depicted by the diagram below.
@@
Therefore, we can interpret (165) as the M5-brane admitting a twisted String structure Sati et al. (2012); Sati (2011), with the twist given by $\mathsf{G}_{4}+\frac{1}{4}\mathsf{p}_{1}]_{\mathbb{Z}}$.
On a related note, based on the arguments in Witten (1998); Freed and Witten (1999) leading to (82), it was conjectured by Witten (2000) that the analogue for the M5-brane should take the form,
@@
For completeness, we expect that (158) should be modified in the presence of sources, analogously to (83). By M/F-theory duality Johnson (1997), the Wilson operators are now M2-branes ending on a 2-submanifold of $X^{6}$. Schematically, the non-vanishing condition for the resulting M5-brane partition function becomes
@@
As alluded to earlier, this relation should be interpreted with care since the M5-brane can no longer be regarded as a decoupled theory with respect to the M-theory background Diaconescu et al. (2003b).
@@ 6 D3-brane in F-theory backgrounds
which is equivalent to $H_{\ast}(T^{2};\mathbb{Z})$ regarded as a $\mathbb{Z}[\pi_{1}(X^{4})]$-module Steenrod (1943); Hatcher (2002).
Notably, in the middle degree we have $\text{Aut}_{\mathbb{Z}}(\mathbb{Z}\oplus\mathbb{Z})=\text{GL}(2,\mathbb{Z})$, i.e. the mapping class group of the torus. For our purposes, it suffices to restrict to the orientation-preserving subgroup $\text{SL}(2,\mathbb{Z})$. As a result, we interpret the resultant D3-brane to be placed in a Type IIB string theory background admitting a non-trivial $\text{SL}(2,\mathbb{Z})$-action. One familiar example of such is an orientifold background, e.g. $X^{4}=S^{1}\times\mathbb{RP}^{3}\subset\text{AdS}_{5}\times\mathbb{RP}^{5}$, where $\rho$ induces a flip in sign of the $\text{SL}(2,\mathbb{Z})$-doublet $([\mathsf{H}_{3}],[\mathsf{G}_{3}])\in H^{3}(\mathbb{RP}^{3};(\mathbb{Z}\oplus\mathbb{Z})_{\rho})$ as we go around a non-contractible loop Witten (1998). [^32] Whenever the context is clear, we will abbreviate $(\mathbb{Z}\oplus\mathbb{Z})_{\rho}$ as $\mathbb{Z}^{2}_{\rho}$. Using the M5-brane basepoint anomaly (157) as the starting point, our goal is to derive the Freed-Witten anomaly cancellation condition for the D3-brane in generic F-theory backgrounds by reducing over the fiber $T^{2}$. [^33] As a reminder, this torus $T^{2}$ is independent from the mapping torus $X^{d}\times S^{1}$.